# Worldline EFT treatment of quadratic and cubic gravity theories

### v1

@@ III Proof of concept: quadratic gravity
$S_{bulk}$ $=$ $\frac{1}{G}\int d^{4}x\hskip 2.84544pt\left[\frac{\phi\nabla^{2}\phi}{8\pi}-\alpha\frac{\nabla^{2}\phi\nabla^{2}\phi}{4\pi}+\frac{1}{c^{2}}\left\{\frac{4(\partial_{t}\phi)^{2}-A_{i}\nabla^{2}A^{i}}{32\pi}+\frac{\alpha}{4\pi}\left(2\nabla^{2}\phi(\partial\phi)^{2}-4\partial^{i}\phi\partial_{i}\partial_{j}\phi\partial^{j}\phi\right.\right.\right.$ (27)
$\left.\left.\left.+2\phi\partial_{i}\nabla^{2}\phi\partial^{i}\phi+6\nabla^{2}\phi(\partial_{t}\phi)^{2}-2\partial^{i}\nabla^{2}\phi\partial_{t}A_{i}-\partial_{i}\nabla^{2}\sigma\partial^{i}\phi-\nabla^{2}\phi\partial_{k}\partial_{j}\sigma^{jk}\right)\right\}+\mathcal{O}\left(\frac{1}{c^{4}}\right)\right]$
where $\sigma=\sigma^{i}_{i}$, $\nabla^{2}\equiv\delta^{ij}\partial_{i}\partial_{j}$, $\delta_{ij}$ is the 3 dimensional Minkowski spatial metric, and $(\partial\phi)^{2}=\partial_{i}\phi\partial^{i}\phi$. We also give the Kaluza-Klein decomposed point particle worldline action for one of the particles till 2 PN for the reader’s convenience
@@ III.1 Newtonian order or 0 PN
In the above diagram, a double line corresponds to a worldline factor, whereas the thin line corresponds to the $\phi$ propagator. The above diagram corresponds to the leading order term in the series expanded generating functional $Z[J]$. The generating functional in quantum field theory encapsulates all correlation functions of a field theory and is defined through a path integral formulation, enabling the systematic computation of n-point functions via functional derivatives with respect to a source term [24]. $Z[J]$ is defined as follows (since the diagram only involves $\phi$ we have suppressed the $A$ and $\sigma$ fields in the following)
@@
III.2 PN1 : Alpha as a perturbing parameter
If we take $\alpha$ as a perturbing parameter, there is no change at 0 PN since the Yukawa potential becomes the Newtonian potential plus a delta function. The delta function will only contribute when the two particles are at the same position, which is not the situation we are looking at. Hence, we can ignore the delta function. Now the interaction terms that we have to consider are at 1 PN and are given by Eq. (27), which leads to the following diagrams with non-zero contribution to the end result
The series expanded generating functional now becomes
$Z$$\approx$ $\frac{1}{2}\int D\phi(k)e^{{-\frac{i}{8\pi G}\int d^{4}k\hskip 2.84544pt\phi(-k)\,\,k^{2}\,\,\phi(k)}}$
$\times$ $\int_{k_{1}...k_{6}}J(-k_{4})J(-k_{5})J(-k_{6})\phi(k_{1})\phi(k_{2})\phi(k_{3})\phi(k_{4})\phi(k_{5})\phi(k_{6})\vec{k_{2}}\cdot\vec{k_{3}}\leavevmode\nobreak\ k_{1}^{2}\delta(\vec{k_{1}}+\vec{k_{2}}+\vec{k_{3}})$
which implies that the amplitude of the $\phi^{3}$ interactions (as in the above diagrams) become
$\mathcal{M}$ $\propto$ $\frac{\alpha}{2}\int_{k_{1}...k_{6}}J(-k_{4})J(-k_{5})J(-k_{6})\left<\phi(k_{1})\phi(k_{2})\phi(k_{3})\phi(k_{4})\phi(k_{5})\phi(k_{6})\right>\vec{k_{2}}\cdot\vec{k_{3}}\leavevmode\nobreak\ k_{1}^{2}\delta(\vec{k_{1}}+\vec{k_{2}}+\vec{k_{3}})$
where $\int_{k_{1}...k_{6}}\equiv\int\prod_{i=1}^{6}\frac{d^{4}k_{i}}{(2\pi)^{4}}$. Performing the integral and using the combinatoric factors, the change in the PN1 result is given as follows
$V_{eff,\alpha}$ $=$ $-\frac{2\alpha G^{2}}{c^{2}}\frac{m_{1}m_{2}(m_{1}+m_{2})}{r_{12}^{4}},$ (42)
where $r_{12}$ is the inter orbital separation. For negative $\alpha$, the result changes by an overall sign.
@@ IV.1 PN-1
At the first PN order, the bulk interaction terms are given as follows:
$\alpha\left[\frac{3\left(\partial_{i}\partial_{i}\phi\right)\left(\partial_{j}\partial_{l}\phi\partial_{j}\partial_{l}\phi\right)}{2G\pi}-\frac{\left(\partial_{i}\partial_{j}\phi\,\partial_{j}\partial_{l}\phi\,\partial_{l}\partial_{i}\phi\right)}{G\pi}\right],$ (45)
@@
Inverse Fourier transforms of the amplitudes of the above two diagrams either vanish or become delta functions with the inter-orbital separation as an argument. Hence, there is no change in the two-body effective potential at the first PN order. We move on to the second PN order.
@@ IV.2 PN-2
In the second PN order, we look at contributions from different Klauza-Klein fields appearing in the bulk interaction. We give the following: the type and kind of fields (or their mixtures) appearing as interaction terms, the interaction terms themselves, and the respective Feynman diagrams when coupled to the worldlines. In the following, double lines denote worldline factors, thin lines denote the $\phi$ propagator, dashed lines denote the $A^{i}$ propagator, and finally, wavy lines denote the $\sigma_{ij}$ propagator. A cross in the middle of a thin line implies that a factor of $\frac{k_{0}}{|\vec{k}|}$ has been inserted in the $\phi$ propagator.
@@
$\frac{-3(\partial_{i}\partial^{i}\phi)^{2}\leavevmode\nobreak\ (\partial_{j}\phi\partial^{j}\phi)}{2G\pi}+\frac{3(\partial_{i}\partial^{i}\phi)\leavevmode\nobreak\ (\partial^{j}\phi\partial_{k}\leavevmode\nobreak\ \partial_{j}\phi\partial^{k}\phi)}{G\pi}+\frac{3(\partial_{i}\phi\partial^{i}\phi)\leavevmode\nobreak\ (\partial_{j}\partial_{k}\phi\leavevmode\nobreak\ \partial^{j}\partial^{k}\phi])}{2G\pi}$
$-\frac{15(\partial_{i}\partial^{j}\phi\leavevmode\nobreak\ \partial^{i}\phi\partial^{k}\phi\leavevmode\nobreak\ \partial_{j}\partial_{k}\phi)}{2G\pi}+\phi\left(\frac{6\partial_{k}\partial^{k}\phi\partial_{i}\partial_{j}\phi\partial^{i}\partial^{j}\phi-4\partial_{i}\partial_{j}\phi\partial^{j}\partial_{k}\phi\partial^{i}\partial^{k}\phi}{G\pi}\right)$ (46)
@@
$-\frac{7(\partial_{i}\partial_{t}\phi\leavevmode\nobreak\ \partial_{j}\partial_{t}\phi\leavevmode\nobreak\ \partial^{i}\phi)}{2G\pi}+\frac{\partial^{i}\partial_{t}\phi\leavevmode\nobreak\ \partial_{j}\partial_{i}\phi\leavevmode\nobreak\ \partial^{j}\partial_{t}\phi}{2G\pi}-3\frac{\partial_{i}\partial_{j}\phi\partial^{i}\partial^{j}\phi\partial_{t}^{2}\phi}{2G\pi}$ (47)
@@
$-\frac{3(\partial_{i}\partial_{t}\phi\leavevmode\nobreak\ \partial^{i}\partial^{j}\phi\leavevmode\nobreak\ \partial_{j}\partial_{k}A^{k}}{4G\pi}-\frac{3(\partial_{i}\partial_{t}A^{j}\leavevmode\nobreak\ \partial^{i}\partial^{k}\phi\leavevmode\nobreak\ \partial_{j}\partial_{k}\phi)}{4G\pi}+\frac{3(\partial_{i}\partial^{j}\phi\leavevmode\nobreak\ \partial^{k}\partial_{k}A^{i}\partial_{j}\partial_{t}\phi)}{4G\pi}-\frac{3(\partial^{j}\partial_{i}A^{i}\partial_{j}\partial^{k}\phi\leavevmode\nobreak\ \partial_{k}\partial_{t}\phi)}{4G\pi}$
$+\frac{3(\partial^{j}\partial_{j}\leavevmode\nobreak\ A^{i}\leavevmode\nobreak\ \partial_{k}\partial_{t}\phi\leavevmode\nobreak\ \partial^{k}\partial_{i}\phi)}{4G\pi}-\frac{3(\partial_{i}\partial_{t}A_{j}\partial^{k}\partial^{i}\phi\leavevmode\nobreak\ \partial^{j}\partial_{k}\phi)}{4G\pi}+\frac{(\partial_{j}\partial_{t}\phi\leavevmode\nobreak\ \partial_{k}\partial_{i}\phi\leavevmode\nobreak\ \partial^{j}\partial^{k}A^{i})}{4G\pi}+\frac{(\partial_{j}\partial_{i}\phi\leavevmode\nobreak\ \partial_{k}\partial_{t}\phi\leavevmode\nobreak\ \partial^{j}\partial^{k}A^{i}}{4G\pi}$
$-\frac{(\partial^{j}\partial^{i}\phi\leavevmode\nobreak\ \partial_{i}\partial_{k}A_{j}\leavevmode\nobreak\ \partial^{k}\partial_{t}\phi)}{4G\pi}-\frac{(\partial^{j}\partial^{i}\phi\leavevmode\nobreak\ \partial_{j}\partial_{k}A_{i}\leavevmode\nobreak\ \partial^{k}\partial_{t}\phi)}{4G\pi}$ (48)

### v2

@@ III Proof of concept: quadratic gravity
$S_{bulk}$ $=$ $\frac{1}{G}\int d^{4}x\hskip 2.84544pt\biggl[-\frac{\left(\partial\phi\right)^{2}}{8\pi}-\frac{\alpha\left(\nabla^{2}\phi\right)^{2}}{4\pi}+\frac{1}{c^{2}}\biggl\{\frac{\partial_{b}A_{a}\partial^{b}A^{a}}{32\pi}+\frac{\left(\partial_{t}\phi\right)^{2}}{8\pi}+\alpha\biggl(-\frac{\phi\left(\nabla^{2}\phi\right)^{2}}{2\pi}$ (27)
$+\nabla^{2}\phi\biggl[\frac{\partial_{t}\partial_{a}A^{a}}{2\pi}+\frac{\left(\partial\phi\right)^{2}}{2\pi}-\frac{\partial_{b}\partial_{a}\sigma^{ab}}{4\pi}+\frac{\nabla^{2}\sigma}{4\pi}+\frac{3\partial_{t}^{2}\phi}{2\pi}\biggr]\biggr)\biggr\}\biggr]$
where $\sigma=\sigma^{i}_{i}$, $\nabla^{2}\equiv\delta^{ij}\partial_{i}\partial_{j}$, $\delta_{ij}$ is the 3 dimensional Minkowski spatial metric, and $(\partial\phi)^{2}=\partial_{i}\phi\partial^{i}\phi$. We also give the Kaluza-Klein decomposed point particle worldline action for one of the particles, till 2 PN, for the reader’s convenience
@@ III.1 Newtonian order or 0 PN
In what follows, we will use double lines denote worldline factors, thin lines to denote the $\phi$ propagator, dashed lines to denote the $A^{i}$ propagator, and finally, wavy lines to denote the $\sigma_{ij}$ propagator. Also, a cross in the middle of a thin line implies that a factor of $\frac{k_{0}}{|\vec{k}|}$ has been inserted in the $\phi$ propagator. The above diagram corresponds to the leading order term in the series expanded generating functional $Z[J]$. The generating functional in quantum field theory encapsulates all correlation functions of a field theory and is defined through a path integral formulation, enabling the systematic computation of n-point functions via functional derivatives with respect to a source term [35]. $Z[J]$ is defined as follows (since the diagram only involves $\phi$ we have suppressed the $A$ and $\sigma$ fields in the following)
@@
However, if the length scale $\propto\sqrt{\alpha}$ is much smaller than the length scales in the potential zone, then the new pole of $R+\alpha R^{2}$ is outside of the cutoff of our approximation and hence the Yukawa like potential goes away. In the following, we show how field redefinition techniques can genuinely get rid of the quadratic corrections to the effective potential.
III.2 $\alpha$ as a perturbing parameter
III.2.1 Till 1 PN
Eq. (27), when treated as a perturbative bulk action in both $\frac{1}{c}$ and $\alpha$, can be simplified further by performing a field redefinition on $\phi$. Considering the following redefinition of $\phi$
$\phi\rightarrow\phi+\frac{2\alpha}{c^{2}}\left(\partial\phi\right)^{2}\,$, (40)
The field redefinition (40) changes the bulk action into the following
$S_{bulk}$ $=$ $\frac{1}{G}\int d^{4}x\hskip 2.84544pt\biggl[-\frac{\left(\partial\phi\right)^{2}}{8\pi}-\frac{\alpha\left(\nabla^{2}\phi\right)^{2}}{4\pi}+\frac{1}{c^{2}}\biggl\{\frac{\partial_{b}A_{a}\partial^{b}A^{a}}{32\pi}+\frac{\left(\partial_{t}\phi\right)^{2}}{8\pi}+\alpha\biggl(-\frac{\phi\left(\nabla^{2}\phi\right)^{2}}{2\pi}$ (41)
$+\nabla^{2}\phi\biggl[\frac{\partial_{t}\partial_{a}A^{a}}{2\pi}-\frac{\partial_{b}\partial_{a}\sigma^{ab}}{4\pi}+\frac{\nabla^{2}\sigma}{4\pi}+\frac{3\partial_{t}^{2}\phi}{2\pi}\biggr]\biggr)\biggr\}\biggr]$
which indicates that the $\nabla^{2}\phi\left(\partial\phi\right)^{2}$ term gets canceled away by the field redefinition. Other $\alpha$ proportional terms lead to contact terms in the effective potential and do not lead to long-range interactions. However, now that the bulk has been modified, the worldline must also change. The worldline till 1 PN gets redefined as follows
$S_{pp}$ $=$ $\int dt\left[-mc^{2}+\frac{1}{2}mv^{2}-m\phi+\frac{m}{c^{2}}\left\{\vec{A}\cdot\vec{v}+\frac{1}{8}|\vec{v}|^{4}-\frac{3}{2}|\vec{v}|^{2}\phi-\frac{1}{2}\phi^{2}-2\alpha\left(\partial\phi\right)^{2}\right\}\right]$
However, it is to be noted that each of the fields appearing in the point particle action is evaluated at the worldline. Therefore, the point particle action can be written as
[…]
$\left.\left.-2\alpha\left(\partial\phi\right)^{2}\right\}\right]$
where $x^{\mu}_{pp}(t)$ is the worldline of one of the point particles. Performing a second field redefinition
$\phi\rightarrow\phi\left[1-8\alpha G\pi m\delta^{4}\left(x^{\mu}-x^{\mu}_{pp}(t)\right)\right]$ (44)
leads to the cancellation of the $\alpha$ dependent term in the worldline against the opposite sign term in the bulk. This has been shown in Appendix A of Ref. [3]. Extra terms coming in the worldline due to the second field redefinition changes the Green’s function only by a overall normalization factor, and the physics remains the same, again as shown in Appendix A of Ref. [3]. Hence, when $\alpha$ is treated as a small parameter, there is no correction to the Einstein-Infeld-Hoffmann (EIH) Lagrangian due to a quadratic correction to the EH action. In the following, we generalize the above method to all PN orders.
III.2.2 For all PN orders
One has the full unperturbed action of the quadratic Ricci scalar correction to the EH action coupled with to a point particle action as
$S_{quad}$ $=$ $\frac{c^{4}}{16\pi G}\int d^{4}x\sqrt{-g}\left(R+\alpha R^{2}\right)-mc^{2}\int d\tau\,d^{4}x\,\delta^{4}\left(x^{\mu}-z^{\mu}(\tau)\right)\sqrt{g_{\mu\nu}u^{\mu}u^{\nu}}$ (45)
Perturbing/varying the action with respect to the metric $g_{\mu\nu}$, that is
$g_{\mu\nu}=\bar{g}_{\mu\nu}+\epsilon h_{\mu\nu}\,$, (46)
for a background $\bar{g}_{\mu\nu}$ and a bookkeeping parameter $\epsilon$, and setting the variation $h_{\mu\nu}$ as
$h_{\mu\nu}=-\frac{\alpha}{\epsilon}\,\,\bar{g}_{\mu\nu}\bar{R}$ (47)
leads to the cancellation of the Ricci scalar squared term in the bulk till $\mathcal{O}\left(\alpha\right)$, and leads to the following change in the net action
$\bar{S}_{quad}$ $=$ $\frac{c^{4}}{16\pi G}\int d^{4}x\sqrt{-\bar{g}}\bar{R}-mc^{2}\int d\tau\,d^{4}x\,\delta^{4}\left(x^{\mu}-z^{\mu}(\tau)\right)\sqrt{\bar{g}_{\mu\nu}u^{\mu}u^{\nu}}$ (48)
$+\frac{1}{2}mc^{2}\alpha\int d\tau\,d^{4}x\,\delta^{4}\left(x^{\mu}-z^{\mu}(\tau)\right)\,\bar{R}\,\sqrt{\bar{g}_{\mu\nu}u^{\mu}u^{\nu}}$
However, this is exactly as one would obtain instead by endowing the point particle action with a non-minimal coupling to the gravitational degrees of freedom in GR. As shown in Appendix A of [3] and in [1], a local field redefinition can eliminate such terms. Hence, when $\alpha$ is treated as a small parameter, there is no change to the effective action at any PN order.
@@ IV.1 PN-1
At the first PN order, the bulk interaction terms for the cubic Riemann part are given as follows:
BIT $=\left[\frac{3\left(\partial_{i}\partial_{i}\phi\right)\left(\partial_{j}\partial_{l}\phi\partial_{j}\partial_{l}\phi\right)}{2G\pi}-\frac{\left(\partial_{i}\partial_{j}\phi\,\partial_{j}\partial_{l}\phi\,\partial_{l}\partial_{i}\phi\right)}{G\pi}\right],$ (51)
@@
The Fourier integral corresponding to the first diagram is given by
$I\propto\int_{k_{1}k_{2}k_{3}}\frac{k_{1}^{2}(k_{2}.k_{3})^{2}}{k_{1}^{2}\penalty 10000\ k_{2}^{2}\penalty 10000\ k_{3}^{2}}\delta^{4}(k_{3}-k_{1}-k_{2})e^{i(k_{1}+k_{2}).x_{1}-ik_{3}.x_{2}}$ (52)
doing integral over $k_{3}$ and defining $p=k_{1}+k_{2}$
$I\propto\int_{k_{2}}\frac{k_{2}^{i}\penalty 10000\ k_{2}^{j}}{k_{2}^{2}}\penalty 10000\ \int_{p}\frac{p^{i}\penalty 10000\ p^{j}}{p^{2}}e^{ip.r_{12}}=0$ (53)
For the second diagram one obtains a vertex factor as $(k_{1}.k_{2})(k_{2}.k_{3})(k_{3}.k_{1})$, which corresponds to the following integral
$I$ $\propto\int_{k_{1}k_{2}k_{3}}\frac{(k_{1}.k_{2})(k_{2}.k_{3})(k_{3}.k_{1})}{k_{1}^{2}\penalty 10000\ k_{2}^{2}\penalty 10000\ k_{3}^{2}}\delta^{4}(k_{3}-k_{1}-k_{2})e^{i(k_{1}+k_{2}).x_{1}-ik_{3}.x_{2}}$ (54)
Integrating over $k_{1}$ one obtains
$I$ $\propto\int_{k_{2}k_{3}}\frac{[(k_{3}-k_{2}).k_{2}](k_{2}.k_{3})[k_{3}.(k_{3}-k_{2})]}{(k_{3}-k_{2})^{2}\penalty 10000\ k_{2}^{2}\penalty 10000\ k_{3}^{2}}e^{ik_{3}.r_{12}}$ (55)
The above corresponds to four terms
$I$ $\propto\Bigg[\int_{k_{2}k_{3}}\frac{(k_{2}.k_{3})^{2}}{(k_{3}-k_{2})^{2}\penalty 10000\ k_{2}^{2}}e^{ik_{3}.r_{12}}-\int_{k_{2}k_{3}}\frac{k_{2}.k_{3}}{(k_{3}-k_{2})^{2}}e^{ik_{3}.r_{12}}$
$-\int_{k_{2}k_{3}}\frac{(k_{2}.k_{3})^{3}}{(k_{3}-k_{2})^{2}\penalty 10000\ k_{2}^{2}\penalty 10000\ k_{3}^{2}}e^{ik_{3}.r_{12}}+\int_{k_{2}k_{3}}\frac{(k_{2}.k_{3})^{2}}{(k_{3}-k_{2})^{2}\penalty 10000\ k_{3}^{2}}e^{ik_{3}.r_{12}}\Bigg]$ (56)
Using the results from the Appendix A of Ref. [37] we find that the integral is $\propto\delta^{\prime\prime\prime}(r_{12})$, where a prime corresponds to a time derivative, and which is zero for non-zero separation of the binary system. Hence, both diagrams evaluate to zero and one has no contribution at the first PN order. We move on to the second PN order.
@@ IV.2 PN-2
In the second PN order, we look at contributions from different Klauza-Klein fields appearing in the bulk interaction. We give the following: the type and kind of fields (or their mixtures) appearing in the interaction, the bulk interaction terms (BIT) themselves, the respective Feynman diagrams when coupled to the worldlines, and some of the calculation details regarding the integrals corresponding to the Feynman diagrams. We have extensively used the results of Appendix A of [37] in calculating the Fourier integrals corresponding to each diagram.
@@
BIT $=\frac{-3(\partial_{i}\partial^{i}\phi)^{2}\penalty 10000\ (\partial_{j}\phi\partial^{j}\phi)}{2G\pi}+\frac{3(\partial_{i}\partial^{i}\phi)\penalty 10000\ (\partial^{j}\phi\partial_{k}\penalty 10000\ \partial_{j}\phi\partial^{k}\phi)}{G\pi}+\frac{3(\partial_{i}\phi\partial^{i}\phi)\penalty 10000\ (\partial_{j}\partial_{k}\phi\penalty 10000\ \partial^{j}\partial^{k}\phi])}{2G\pi}$
$-\frac{15(\partial_{i}\partial^{j}\phi\penalty 10000\ \partial^{i}\phi\partial^{k}\phi\penalty 10000\ \partial_{j}\partial_{k}\phi)}{2G\pi}+\phi\left(\frac{6\partial_{k}\partial^{k}\phi\partial_{i}\partial_{j}\phi\partial^{i}\partial^{j}\phi-4\partial_{i}\partial_{j}\phi\partial^{j}\partial_{k}\phi\partial^{i}\partial^{k}\phi}{G\pi}\right)$ (57)
A typical vertex factor for the $\phi^{4}$ interaction arises from the four $\phi$ interaction term in the Lagrangian. All momenta involved in the interaction are 3-vectors except for the delta function, which ensures momentum conservation. By integrating over $k_{1}$ using the delta function, we eliminate one of the momenta. Subsequently, performing the remaining integrals yields a result that is proportional to $1/r_{12}^{7}$.
The expression for a typical vertex factor, along with the propagator and momentum conservation, is given by:
$\mathcal{V}=\frac{k_{1}^{2}(k_{2}\cdot k_{3})(k_{3}\cdot k_{4})}{k_{1}^{2}k_{2}^{2}k_{3}^{2}k_{4}^{2}}\delta^{4}(k_{1}+k_{2}+k_{3}+k_{4})$ (58)
Hence, a typical integrand looks like
$\mathcal{I}=\frac{k_{1}^{2}(k_{2}\cdot k_{3})(k_{3}\cdot k_{4})}{k_{1}^{2}k_{2}^{2}k_{3}^{2}k_{4}^{2}}\delta^{4}(k_{1}+k_{2}+k_{3}+k_{4})\times e^{i(k_{1}\cdot x_{1})}e^{[i(k_{2}+k_{3}+k_{4})\cdot x_{2}]}$ (59)
where $k_{1},k_{2},k_{3},k_{4}$ are the 3-momenta of the vertex factor, the delta function $\delta^{4}(k_{1}+k_{2}+k_{3}+k_{4})$ enforces four-momentum conservation at the vertex, and the denominator accounts for the propagators of the intermediate particles.
After summing over all possible scalar product combinations corresponding to the $\phi^{4}$ diagrams, the final result takes the form:
$I=\frac{160G^{3}\alpha(m_{1}m_{2}^{3}+m_{1}^{3}m_{2})}{7r_{12}^{7}}$ (60)
@@
BIT $=-\frac{7(\partial_{i}\partial_{t}\phi\penalty 10000\ \partial_{j}\partial_{t}\phi\penalty 10000\ \partial^{i}\phi)}{2G\pi}+\frac{\partial^{i}\partial_{t}\phi\penalty 10000\ \partial_{j}\partial_{i}\phi\penalty 10000\ \partial^{j}\partial_{t}\phi}{2G\pi}-3\frac{\partial_{i}\partial_{j}\phi\partial^{i}\partial^{j}\phi\partial_{t}^{2}\phi}{2G\pi}$ (61)
A typical vertex factor for the $\phi^{3}$ interaction terms comes with a time derivative, which, along with the propagator, is given by:
$\mathcal{V}=\frac{(k_{1}\cdot k_{2})^{2}\omega_{3}^{2}}{k_{1}^{2}k_{2}^{2}k_{3}^{2}}\delta^{4}(k_{1}+k_{2}+k_{3}).$ (62)
The $\omega_{3}$ comes from the time derivative acting on the $e^{i\omega_{3}t}$, which itself comes due to $\phi$ being written in the Fourier space, as in Eq. (29), where $k_{0}=\frac{\omega}{c}$. The above gravitational interaction leads to the following contributions to the effective Lagrangian:
$I$ $=\frac{G^{2}m_{1}m_{2}^{2}}{c^{4}}\bigg[\frac{1}{r_{12}^{6}}\bigg(12\alpha(v_{1}^{i}v_{1}^{i})+24\alpha(v_{1}^{i}v_{2}^{i})+12\alpha(v_{2}^{i}v_{2}^{i})\bigg)$
$+\frac{1}{r_{12}^{8}}\bigg(-72\alpha(v_{1}^{i}x^{i})^{2}-144\alpha(v_{1}^{i}x^{i})(v_{2}^{i}x^{i})-72\alpha(v_{2}^{i}x^{i})^{2}\bigg)$
$+\frac{1}{r_{12}^{6}}\bigg(12\alpha(a_{1}^{i}x^{i})-12\alpha(a_{2}^{i}x^{i})\bigg)\bigg]$
$+\frac{G^{2}m_{1}^{2}m_{2}}{c^{4}}\bigg[\frac{1}{r_{12}^{6}}\bigg(12\alpha(v_{1}^{i}v_{1}^{i})+24\alpha(v_{1}^{i}v_{2}^{i})+12\alpha(v_{2}^{i}v_{2}^{i})\bigg)$
$+\frac{1}{r_{12}^{8}}\bigg(-72\alpha(v_{1}^{i}x^{i})^{2}-144\alpha(v_{1}^{i}x^{i})(v_{2}^{i}x^{i})-72\alpha(v_{2}^{i}x^{i})^{2}\bigg)$
$+\frac{1}{r_{12}^{6}}\bigg(12\alpha(a_{1}^{i}x^{i})-12\alpha(a_{2}^{i}x^{i})\bigg)\bigg].$ (63)
This result shows how the gravitational interaction modifies the dynamics at different orders, leading to corrections at $r_{12}^{-5}$ and beyond. The $\omega_{3}^{2}$ leads to a velocity or an acceleration term in the results. $x^{i}$ is the radial vector connecting $m_{1}$ to $m_{2}$, $v_{1}^{i},v_{2}^{i}$ are the velocity 3-vectors, and $a_{1}^{i},a_{2}^{i}$ represent acceleration 3-vectors of the two particles.
@@
BIT $=-\frac{3(\partial_{i}\partial_{t}\phi\penalty 10000\ \partial^{i}\partial^{j}\phi\penalty 10000\ \partial_{j}\partial_{k}A^{k}}{4G\pi}-\frac{3(\partial_{i}\partial_{t}A^{j}\penalty 10000\ \partial^{i}\partial^{k}\phi\penalty 10000\ \partial_{j}\partial_{k}\phi)}{4G\pi}+\frac{3(\partial_{i}\partial^{j}\phi\penalty 10000\ \partial^{k}\partial_{k}A^{i}\partial_{j}\partial_{t}\phi)}{4G\pi}$
$-\frac{3(\partial^{j}\partial_{i}A^{i}\partial_{j}\partial^{k}\phi\penalty 10000\ \partial_{k}\partial_{t}\phi)}{4G\pi}+\frac{3(\partial^{j}\partial_{j}\penalty 10000\ A^{i}\penalty 10000\ \partial_{k}\partial_{t}\phi\penalty 10000\ \partial^{k}\partial_{i}\phi)}{4G\pi}-\frac{3(\partial_{i}\partial_{t}A_{j}\partial^{k}\partial^{i}\phi\penalty 10000\ \partial^{j}\partial_{k}\phi)}{4G\pi}$
$+\frac{(\partial_{j}\partial_{t}\phi\penalty 10000\ \partial_{k}\partial_{i}\phi\penalty 10000\ \partial^{j}\partial^{k}A^{i})}{4G\pi}+\frac{(\partial_{j}\partial_{i}\phi\penalty 10000\ \partial_{k}\partial_{t}\phi\penalty 10000\ \partial^{j}\partial^{k}A^{i}}{4G\pi}-\frac{(\partial^{j}\partial^{i}\phi\penalty 10000\ \partial_{i}\partial_{k}A_{j}\penalty 10000\ \partial^{k}\partial_{t}\phi)}{4G\pi}$
$-\frac{(\partial^{j}\partial^{i}\phi\penalty 10000\ \partial_{j}\partial_{k}A_{i}\penalty 10000\ \partial^{k}\partial_{t}\phi)}{4G\pi}$ (64)
The typical vertex factor along with the propagator and momentum conservation is given as follows:
$\mathcal{V}=\frac{k_{1}\cdot v_{j}\,k_{2}^{2}\,k_{1}\cdot k_{3}\,\omega_{3}}{k_{1}^{2}k_{2}^{2}k_{3}^{2}}\delta^{4}(k_{1}+k_{2}+k_{3})\qquad\qquad j=1,2$ (65)
It contains one time derivative but also includes terms of order $v_{j}^{i}$. As a result, this interaction contributes to either velocity-squared or acceleration dependent terms. The gravitational interaction at this order leads to the following contribution to the effective Lagrangian:
$I$ $=\frac{G^{2}}{c^{4}}\bigg[m_{1}^{2}m_{2}\bigg(\frac{-24v_{2}^{2}-72(v_{1}^{i}v_{2}^{i})+24(a_{2}^{i}x^{i})-72(a_{1}^{i}x^{i})}{r_{12}^{6}}+\frac{216(v_{1}^{i}x^{i})(v_{2}^{i}x^{i})+144(v_{2}^{i}x^{i})^{2}+216(v_{1}^{i}x^{i})^{2}}{r_{12}^{8}}\bigg)$
$+m_{1}m_{2}^{2}\bigg(\frac{-24v_{1}^{2}-72(v_{1}^{i}v_{2}^{i})+72(a_{2}^{i}x^{i})-24(a_{1}^{i}x^{i})}{r_{12}^{6}}\quad+\frac{216(v_{1}^{i}x^{i})(v_{2}^{i}x^{i})+216(v_{2}^{i}x^{i})^{2}+144(v_{1}^{i}x^{i})^{2}}{r_{12}^{8}}\bigg)\bigg]$